# ]Meta FAIR Improving Factuality with Explicit Working Memory

### v1

@@ Abstract
Large language models can generate factually inaccurate content, a problem known as hallucination. Recent works have built upon retrieved-augmented generation to improve factuality through iterative prompting but these methods are limited by the traditional RAG design. To address these challenges, we introduce Ewe (Explicit Working Memory), a novel approach that enhances factuality in long-form text generation by integrating a working memory that receives real-time feedback from external resources. The memory is refreshed based on online fact-checking and retrieval feedback, allowing Ewe to rectify false claims during the generation process and ensure more accurate and reliable outputs. Our experiments demonstrate that Ewe outperforms strong baselines on four fact-seeking long-form generation datasets, increasing the factuality metric, VeriScore, by 2 to 10 points absolute without sacrificing the helpfulness of the responses. Further analysis reveals that the design of rules for memory updates, configurations of memory units, and the quality of the retrieval datastore are crucial factors for influencing model performance.
@@ 1 Introduction
We demonstrate empirically that Ewe generates more factual responses without sacrificing the relevance to the input questions, using four fact-seeking long-form generation datasets. In general, with the feedback from online fact-checking and targeted retrieval, Ewe increases VeriScore (Song et al., 2024), the factuality metric we use, by 2 to 10 points absolute and has a similar helpfulness in terms of instruction following compared to the base model Llama-3.1.
@@ 4.3 Baselines
We used instruction-tuned Llama-3.1 70B and 8B as the base models and compared Ewe with five baselines: base model only, retrieval augmentation (RA), Chain of verification (CoVe)555We adapted an implementation from https://github.com/ritun16/chain-of-verification, an iterative retrieval approach DRAGIN (Su et al., 2024)666We used the authors’ implementation https://github.com/oneal2000/DRAGIN, and a recently proposed semi-parametric decoding method Nest (Li et al., 2024). For base model only, Llama-3.1 or Llama-3.1, we simply gave the language model the prompt in the dataset and the instruction of requesting detailed information, without other additional information. With retrieval augmentation, we retrieved 20 passages using the input prompts as queries and then prepended the passages to the input777Using more than 20 passages does not provide significant benefits in our preliminary experiments, so we limit our retrieval to the top 20 passages.. Nest is a strong retrieval-based decoding algorithm. Following the original setup, we retrieved 100 passages to use as candidates. For CoVe, we employ the “factor+revise” method, which Dhuliawala et al. (2024) demonstrated to be the most effective. Additionally, we improve CoVe by integrating retrieved passages from our retrieval datastore during the verification step. This augmentation helps us establish a stronger and more comparable baseline method, considering that most other baseline methods also utilize retrieval. For all our experiments, the maximum generation step was set to 1024. Llama-3.1 is used as the baseline method for all AlpacaEval comparisons.
@@ 5.1 Memory Configurations
In this analysis, we explore the influence of memory configurations on factuality, based on experiments on 50 randomly sampled prompts from LongFact. We examine this impact through two dimensions: the number of memory units and the shape of memory units.
In Figure 3, we investigate how varying the numbers of memory units used for storing fact-checking feedback and retrieved passages may impact factuality. When adjusting the configuration for one, we keep the other constant to facilitate easier interpretation of the results. Overall, we observe that having a large amount of memory units for either fact-checking feedback or retrieved passages negatively impacts factuality. This is likely because a significant amount of stale information remains in working memory for an extended period without being updated, as we adhere to the FIFO rule for updating working memory. Consequently, this information becomes outdated as the generation process continues.
| Memory Shape () | |||
| F1 | 74.0 | 69.0 | 68.2 |
In Table 2, we examine the impact of varying memory unit shapes on factuality. To ensure a fair comparison, we maintained a consistent total number of tokens in working memory across different experimental setups. Notably, our findings suggest that models favor shorter, more memory units over longer, fewer ones. We hypothesize that this preference arises because 128 tokens approximately match the length of a retrieved passage, allowing the attention mechanism to effectively cover one individual passage at a time. In contrast, longer memory units combine multiple passages into a single unit, which may compel the attention to focus on less relevant passages when they are grouped with more relevant ones.
@@ 5.3 Model Confidence
As shown in Figure 4(a), when using a fixed retrieval interval, an intermediate interval seems to perform well. This may be due to the fact that overly frequent retrieval can add irrelevant and conflicting information to the memory. With fact-checking feedback in memory, it seems frequent verification and regeneration is not always beneficial due to the fact that we only regenerate once and the regenerated sentence is not always better.
@@
In practice, fixed retrieval and verification intervals may be unnecessary and lead to sub-optimal performance. We explore whether model-confidence can serve as a signal for refreshing working memory. Specifically, we compare two different metrics for model confidence: (1) Entropy: average entropy of generated tokens in a sentence, and (2) Min-prob: minimum probability of tokens in a sentence. A higher threshold for entropy results in less frequent memory update, and a higher threshold for min-prob results in more frequent memory update. Since external fact-checkers can be computationally expensive, we first examine if we can use model confidence as a signal for retrieval and regeneration, without using an auxiliary fact-checking model to provide feedback. As shown in Figure 4(b) and 4(c) (blue line), we observe empirically intermediate thresholds for retrieval perform well, leading to to better F1 when compared to the settings in Figure 4(a), where we use different fixed intervals for retrieval. With external fact-checkers, we investigate if we can use model confidence as a signal to trigger verification and regeneration to improve generation efficiency. In Figure 4(b) and 4(c) (red line), when chosen at an appropriate threshold, both entropy and min-prob can outperform the baseline (using fixed verification interval ) despite with less frequent verification.
@@ 5.4 Knowledge from Retrieval
| Wiki | 63.0 | 39.1 | 55.4 | 50.5 |
| C4 | 66.8 | 45.4 | 58.6 | 53.6 |
| C4 + Wiki | 69.0 | 43.8 | 58.8 | 54.6 |
We present the results of using different retrieval corpora in Table 4, including Wikipedia, C4, or both of them together. Likely due to its broader coverage, C4 is more effective than Wikipedia in helping the model produce more factual responses. Combining C4 with Wikipedia further enhances the factual accuracy (except for Biography), probably because they offer complementary sets of knowledge.
@@ Acknowledgments
We thank Victoria Lin and Barlas Oğuz for their insightful feedback, and we are grateful to Maria Lomeli for her support in building the retrieval system.

### v2

@@ Abstract
Large language models can generate factually inaccurate content, a problem known as hallucination. Recent works have built upon retrieved-augmented generation to improve factuality through iterative prompting but these methods are limited by the traditional RAG design. To address these challenges, we introduce Ewe (Explicit Working Memory), a novel approach that enhances factuality in long-form text generation by integrating a working memory that receives real-time feedback from external resources. The memory is refreshed based on online fact-checking and retrieval feedback, allowing Ewe to rectify false claims during the generation process and ensure more accurate and reliable outputs. Our experiments demonstrate that Ewe outperforms strong baselines on four fact-seeking long-form generation datasets, increasing the factuality metric, VeriScore, by 2 to 6 points absolute without sacrificing the helpfulness of the responses. Further analysis reveals that the design of rules for memory updates, configurations of memory units, and the quality of the retrieval datastore are crucial factors for influencing model performance.
@@ 1 Introduction
We demonstrate empirically that Ewe generates more factual responses without sacrificing the relevance to the input questions, using four fact-seeking long-form generation datasets. In general, with the feedback from online fact-checking and targeted retrieval, Ewe increases VeriScore (Song et al., 2024), the factuality metric we use, by 2 to 6 points absolute and has a similar helpfulness in terms of instruction following compared to the base model Llama-3.1. We also perform human evaluations to ensure that our enhancements in VeriScore effectively translate to factuality.
@@ 4.3 Baselines
We used instruction-tuned Llama-3.1 70B and 8B as the base models and compared Ewe with five baselines: base model only, retrieval augmentation (RA), Chain of verification (CoVe)555We adapted an implementation from https://github.com/ritun16/chain-of-verification, an iterative retrieval approach DRAGIN (Su et al., 2024)666We used the authors’ implementation https://github.com/oneal2000/DRAGIN, and a recently proposed semi-parametric decoding method Nest (Li et al., 2024). For base model only, Llama-3.1 or Llama-3.1, we simply gave the language model the prompt in the dataset and the instruction of requesting detailed information, without other additional information. With retrieval augmentation, we retrieved 10 passages using the input prompts as queries and then prepended the passages to the input777Using more than 10 passages does not provide significant benefits in our preliminary experiments, so we limit our retrieval to the top 10 passages.. Nest is a strong retrieval-based decoding algorithm. Following the original setup, we retrieved 100 passages to use as candidates. For CoVe, we employ the “factor+revise” method, which Dhuliawala et al. (2024) demonstrated to be the most effective. Additionally, we improve CoVe by integrating retrieved passages from our retrieval datastore during the verification step. This augmentation helps us establish a stronger and more comparable baseline method, considering that most other baseline methods also utilize retrieval. For all our experiments, the maximum generation step was set to 1024. Llama-3.1 is used as the baseline method for all AlpacaEval comparisons.
@@ 5.1 Memory Configurations
In this analysis, we explore the influence of memory configurations on factuality, based on experiments on 50 randomly sampled prompts from LongFact. We examine this impact through two dimensions: the number of memory units and the lengths of memory units.
In Figure 3, we investigate how varying the numbers of memory units used for storing fact-checking feedback and retrieved passages may impact factuality. We apply the same configuration to both, as similar trends are observed when adjusting the settings for each type of memory. Overall, we find that having a large number of memory units for either fact-checking feedback or retrieved passages negatively affects precision and, consequently, factuality (though recall remains largely unaffected). This is likely because a significant amount of stale information remains in working memory for an extended period without being updated, as we adhere to the FIFO rule for updating working memory. Consequently, this information becomes outdated as the generation process continues.
In Figure 4, we examine the impact of varying lengths of memory units on factuality. Our results indicate that shorter memory units tend to increase precision but decrease recall, whereas longer memory units tend to decrease precision but increase recall. We hypothesize that this occurs because short memory units allow the attention mechanism to allocate weights more effectively to individual passages. In contrast, longer memory units combine multiple passages into a single unit, which may compel the attention to focus on less relevant passages when they are grouped with more relevant ones. However, longer memory units also provide advantages in terms of improved reasoning across documents, which appear to be beneficial for recall.
In Figure 5, we examine the impact of varying retrieval score thresholds. Notably, initially raising the threshold (allowing more relevant information to be retained in working memory) enhances both precision and recall. Further increasing the threshold continues to improve precision while maintaining recall at a similar level. However, increasing the threshold beyond this point negatively affects F1 and recall, despite still benefiting precision, likely due to the exclusion of too many retrievals. These findings indicate that our model might gain from a more advanced memory updating mechanism, which we intend to investigate in future work.
@@ 5.3 Model Confidence
As shown in Figure 6(a), when using a fixed retrieval interval, an smaller interval seems to perform better in general with a few spikes. This may be due to the fact that overly frequent retrieval can add irrelevant and conflicting information to the memory. On the other hand, when using a fixed verification interval, shorter interval always leads to better performance.
@@
In practice, fixed retrieval and verification intervals may be unnecessary and lead to sub-optimal performance. We explore whether model-confidence can serve as a signal for refreshing working memory. Specifically, we compare two different metrics for model confidence: (1) Entropy: average entropy of generated tokens in a sentence, and (2) Min-prob: minimum probability of tokens in a sentence. A higher threshold for entropy results in less frequent memory update, and a higher threshold for min-prob results in more frequent memory update. Since external fact-checkers can be computationally expensive, we first examine if we can use model confidence as a signal for retrieval and regeneration, without using an auxiliary fact-checking model to provide feedback. As shown in Figure 6(b) and 6(c) (blue line), we observe empirically intermediate thresholds for retrieval perform well, leading to to better F1 when compared to the settings in Figure 6(a), where we use different fixed intervals for retrieval. With external fact-checkers, we investigate if we can use model confidence as a signal to trigger verification and regeneration to improve generation efficiency. In Figure 6(b) and 6(c) (green line), when chosen at an appropriate threshold, both entropy and min-prob can outperform the baseline (using fixed verification interval with the same memory configuration) despite with less frequent verification.
@@ 5.4 Knowledge from Retrieval
| Wiki | 67.9 | 46.1 | 55.5 | 52.5 |
| C4 | 70.8 | 44.6 | 53.7 | 53.3 |
| C4 + Wiki | 74.8 | 48.4 | 53.3 | 52.3 |
We present the results of using different retrieval corpora in Table 3, including Wikipedia, C4, or both of them together. Interestingly, different datasets exhibit distinct preferences. For LongFact and Fava, C4 proves more effective than Wikipedia in helping the model to generate more factual responses. Conversely, Biography and AlpacaFact show a preference for Wikipedia. Combining C4 with Wikipedia further improves factual accuracy for LongFact and Biography, but the same trend is not observed for AlpacaFact and Fava. This is likely because the first two datasets consist of domains where C4 and Wikipedia provide complementary knowledge, whereas the latter two do not.
In Figure 7, we explore the impact of using different sizes of retrieval datastores. Both LongFact and Biography exhibit similar patterns, where increasing the datastore size generally improves results. In contrast, Fava and AlpacaFact differ, as Ewe tends to maintain similar level of VeriScore F1 across various setups. We hypothesize that real-time feedback from fact verifiers helps offset the lack of information from retrieval datastores in these two datasets.
5.5 Human Evaluation
| RA | 0.61 |
| Ewe | 0.65 |
We conduct a human evaluation to determine if using VeriScore during inference impacts its alignment with human judgments. To achieve this, we randomly select 100 generation samples each from our retrieval-augmented baseline and Ewe, and recruit 6 human annotators. During sample selection, we categorize sentences into buckets based on the number of nonfactual claims predicted by VeriScore. We then randomly choose samples from each bucket to ensure an equal representation of sentences with varying numbers of nonfactual claims. Annotators are given a sentence from model responses along with the full response for context. They are instructed to classify the sentence as factual, nonfactual, or “unable to decide” if no claim is present. Additionally, they are encouraged to use online search engines if they are unsure about specific information. Our evaluation gathers 120 annotations. Since annotators might choose “unable to decide” for various reasons, such as the sentence being functional and not making any claims, we exclude these from score computation. We calculate Cohen’s Kappa to assess inter-annotator agreement between humans and VeriScore, excluding 25 “unable to decide” annotations. The results are presented in Table 4. Since a Cohen’s Kappa score above 0.61 is considered substantial and the scores do not decrease with Ewe, we conclude that using VeriScore during inference has indeed enhanced the factuality of Ewe.
@@ Acknowledgments
We thank Victoria Lin, Barlas Oğuz, and Xilun Chen for their insightful feedback, and we are grateful to Maria Lomeli for her support in building the retrieval system.